OpenStudy (anonymous):

Which equation could be used to solve the question below? Tirza has b books. Her mom has 36 books and her sister has 23 books. They have 82 books in all. How many books does Tirza have? b - 36 + 23 = 82 b + 36 - 23 = 82 b + 36 + 23 = 82 b - 36 - 23 = 82

OpenStudy (anonymous):

Fan and Medal!!!!!

OpenStudy (anonymous):

Which equation could be used to solve the question below? Kitty made b cups of dough. She divided the dough into 3 equal clumps. Each of the dough clumps doubled in size. Now, each clump is made of 6 cups of dough. How many cups of dough did Kitty originally make? b × 3 × 2 = 6 b ÷ 3 × 2 = 6 b × 6 ÷ 2 = 6 b ÷ 3 ÷ 2 = 6

OpenStudy (texaschic101):

b + 36 + 23 = 82 <--- your equation b + 59 = 82 b = 82 - 59 b = 23

OpenStudy (texaschic101):

b/3 * 2 = 6